# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models

### v2

@@ Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect. Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement and analyze an orthonormal regularizer, which could effectively transfer the scaling information into $\Sigma_{p}$ and ultimately allows the training process to be more efficient. SORSA adapters could be merged during inference, thus eliminating any inference latency. After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments. On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%). We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
Deep Learning, LLM, PEFT, LoRA, SVD
@@ 2 Related Works
Adapter-based PEFT is the first type of PEFT, which was initially designed by (Houlsby et al., 2019). It introduces additional trainable non-linear blocks into the frozen pre-trained model, which could effectively tune the pre-trained model with a limited amount of trainable parameters. Its variants, e.g., (Lin et al., 2020), reduce the number of adapter layers per block, and (He et al., 2022) focus on adding adapter modules parallel to existing layers. However, all adapter-based PEFT methods introduce inference latency due to their non-mergeable attribute.
Prompt-based PEFT is also a well-known type of PEFT, which was first proposed in (Lester et al., 2021). There are several variants of this work, including (Liu et al., 2022; Razdaibiedina et al., 2023). However, they have some inevitable shortcomings, such as potential performance limitations compared to full parameter fine-tuned models, additional inference latency due to expanding the length of the total input to the model, and the complexity of designing effective initialization.
@@ 3.3 Analysis Result
4 Our Method
4.1 Singular Value and Orthonormal Regularized Singular Vector Adaptation
@@
The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024). Nevertheless, unlike PiSSA which merge $\Sigma_{p}$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\Sigma_{p}^{\frac{1}{2}}$ and $B=\Sigma_{p}^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ in separate matrices. SORSA is defined by Eq (8), which is initially equivalent to the pre-trained weight $W_{0}$.
@@
We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 4.2.
$\mathcal{L}_{reg}=\|U^{\top}_{p}U_{p}-I\|_{F}+\|V_{p}V^{\top}_{p}-I\|_{F}$ (9)
@@
At training step $t$, $\nabla_{W_{t}}\mathcal{L}_{train}$ denotes the gradient of $W_{t}$ respect to $\mathcal{L}_{train}$, and $\nabla_{W_{t}}\mathcal{L}_{reg}$ denotes the gradient of $W_{t}$ respect to the orthonormal regularizer loss $\mathcal{L}_{reg}$. $\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss, respectively.
We update the SORSA as Eq (11) for implementation simplicity.
@@
$\eta_{d}$ is the learning rate which is given to the scheduler. This implementation allows us to use only one optimizer and scheduler to deal with two different learning rates separately.
4.2 Gradient Analysis
In this section, we provides a gradient analysis for the SORSA with the orthonormal regularizer, focusing on proving the regularizer’s convexity and find the upper bound of the hyperparameter $\gamma$. We also demonstrates how the regularizer stabilizes the training and facilitates its generalization.
To prove SORSA with regularizer is optimizable, we need to prove the convexity of the regularizer, and then use Lipschitz continuity to find the upper bound of $\gamma$.
In order to analyze the regularizer according to Eq. (9), we begin by noticing that the Frobenius norm $\|\cdot\|_{F}$ is a convex function. We then examine the terms $U^{\top}_{p}U_{p}-I$ and $V_{p}V^{\top}_{p}-I$, which are quadratic forms in $U_{p}$ and $V_{p}$ , respectively. To assess their convexity, we compute the Hessian matrix of each quadratic term with respect to $U_{p}$ and $V_{p}$. For the term $U^{\top}_{p}U_{p}-I$ , we first express it as a function of $U_{p}$. The Hessian of a quadratic form $U^{\top}_{p}U_{p}$ is given by $2I$, where $I$ is the identity matrix of appropriate dimensions. This is because the second-order derivative of $U^{\top}_{p}U_{p}$ with respect to $U_{p}$ is constant and positive semidefinite, ensuring convexity. The second term could use the exact same process, which is also convex. Therefore, the regularizer is convex.
Then, we analyze the Lipschitz continuity of the gradients to establish an appropriate bound for $\gamma$. A function $f$ is Lipschitz continuous with constant $L$ if:
$\|\nabla f(W_{1})-\nabla f(W_{2})\|\leq L\|W_{1}-W_{2}\|$ (12)
According to Eq. (11), the gradient of combined loss function $\mathcal{L}(W_{p})$ is:
$\nabla_{W_{p,t}}\mathcal{L}=\nabla_{W_{p,t}}\mathcal{L}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}\mathcal{L}_{reg}$ (13)
Assuming that $\nabla_{W_{p}}\mathcal{L}_{train}$ and $\nabla_{W_{p}}\mathcal{L}_{reg}$ are Lipschitz continuous with constants $L_{train}$ and $L_{reg}$, respectively, the Lipschitz constant for the combined loss function can be written as:
$L=L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}$ (14)
When the learning rate is at maximum point, $\eta_{t}=\eta_{d}$, to ensure convergence of the algorithm, the learning rate $\eta_{d}$ must satisfy the following condition:
$\eta_{d}<\frac{2}{L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}}$ (15)
Which could be written as:
$\gamma<\frac{2-\eta_{d}L_{train}}{L_{reg}}$ (16)
$\gamma<\frac{2}{k\eta_{d}}-\frac{L_{train}}{k}$ (17)
This inequality provides an upper bound on $\gamma$ to ensure that the regularizer does not prevent convergence when $\eta_{t}$ is at its maximum value $\eta_{d}$. The inequality shows that $\gamma$ should have inversely proportional relationship to $\eta_{d}$ while selecting the hyperparameter in order to make sure the SORSA adapters could converge.
Figure 3: Orthonormality, loss and gradient norm variation of SORSA using different $\gamma$ on MetaMathQA training of Mistral 7B v0.1.
In Figure 3, we could see that below the bound, $\gamma$ have almost no effect on loss curve, reflecting the robustness of SORSA.
The reinforcement of orthonormality of $U_{p}$ and $V^{\top}_{p}$ concentrates the scaling information into $\Sigma_{p}$, leads to a more holistic parameter updating process. The regularizer ensures that $U_{p}$ and $V^{\top}_{p}$ act as pure rotation matrices, while all magnitude updates are captured within $\Sigma_{p}$ by decoupling singular values from singular vectors,
Moreover, training along with an Adam-like (Kingma & Ba, 2015) optimizer could accelerate this process because its momentum on parameters of $U_{t}$ will become much smaller than momentum on parameters of $\Sigma_{t}$, which means the optimizer will halt the parameter updating towards the direction point to orthonormality increasing in $U_{p}$ and $V^{\top}_{p}$, and more likely to update the scaling information to $\Sigma_{p}$.
This separation stabilizes learning, as the singular vectors remain orthonormal, preventing distortions and over-fitting. Furthermore, concentrating the updates within $\Sigma_{p}$ allows for a more efficient gradient flow, facilitating smoother updates and ensuring that scaling changes reflect true data variation. Consequently, this structured update mechanism fosters better generalization and balanced learning throughout the training process.
Introducing the orthonormal regularizer into the training process promotes a more stable update trajectory and a more evenly distributed weight by actively countering deviations from orthonormality. This regulation stabilizes the updates and enhances the model’s ability to maintain crucial properties of the singular vectors $U_{p}$ and $V^{\top}_{p}$, thereby contributing to improved generalization and effectiveness in learning. This analysis underlines the utility of orthonormal regularization in complex model training scenarios, particularly in fine-tuning pre-trained models where preserving learned features is crucial.
@@ 5 Empirical Experiments
We conducted NLG tests on Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023). We trained the models with SORSA applied and conducted GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) tests. The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024). See Appendix A.2 for more details and hyperparameters of the training. We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from the PiSSA paper (Meng et al., 2024). For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80G) GPU. See Table 1 for the results and Figure 4 for the loss and gradient norm comparison. See Figure 5 for the variation in orthonormality of SORSA.
Model Method Trainable Parameters MATH GSM-8K
Llama 2 7B Full FT 6738M 7.22 49.05
LoRA 320M 5.50 42.30
PiSSA 320M 7.44 53.07
SORSA 320M 10.36 56.03
Mistral 7B Full FT 7242M 18.60 67.02
LoRA 168M 19.68 67.70
PiSSA 168M 21.54 72.86
SORSA 168M 21.86 73.09
Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from the paper of PiSSA (Meng et al., 2024).
@@
Figure 4: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B.
Figure 5: Variation of orthonormality on MetaMathQA training of Llama 2 7B and Mistral 7B v0.1.
@@ 6 Conclusion
We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis. We also analyze the regularizer of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
@@ 7 Future Work
  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different transformer-based LLMs. This includes replicating other tests from studies like PiSSA (Meng et al., 2024) and performing evaluations such as HumanEval (Chen et al., 2021) and MT-Bench (Zheng et al., 2023), in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
  • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2021a) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.
  • Application to Other Fields: Test SORSA on different fields beyond NLP. Such as conducting experiments on computer vision (CV) models like DDPM (Ho et al., 2020), DDIM (Song et al., 2020), and ViT (Dosovitskiy et al., 2021). These experiments could expand SORSA’s applicability to a broader range of deep-learning tasks.

### v3

@@ Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect. Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement and analyze an orthonormal regularizer, which we prove could decrease the condition number of $W_{p}$ and allows the optimization to be more efficient. SORSA adapters could be merged during inference, thus eliminating any inference latency. After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments. On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%). We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
@@ 1 Introduction
We also provide a comprehensive gradient analysis that provides a mathematical foundation for SORSA. This analysis demonstrates several crucial properties of our method, including the convexity of the regularizer, Lipschitz continuity of the gradient, bounds on the hyperparameter $\gamma$. Moreover, we prove that SORSA improves the condition number of the optimization problem compared to unregularized approaches.
@@ 2 Related Works
Adapter-based PEFT is the first type of PEFT, which was initially designed by Houlsby et al. (2019). It introduces additional trainable non-linear blocks into the frozen pre-trained model, which could effectively tune the pre-trained model with a limited amount of trainable parameters. Its variants, e.g., Lin et al. (2020), reduce the number of adapter layers per block, and He et al. (2022) focus on adding adapter modules parallel to existing layers. However, all adapter-based PEFT methods introduce inference latency due to their non-mergeable attribute.
Prompt-based PEFT is also a well-known type of PEFT, which was first proposed in Lester et al. (2021). There are several variants of this work, including Liu et al. (2022); Razdaibiedina et al. (2023). However, they have some inevitable shortcomings, such as potential performance limitations compared to full parameter fine-tuned models, additional inference latency due to expanding the length of the total input to the model, and the complexity of designing effective initialization.
@@ 3.3 Analysis Result
4 Singular Value and Orthonormal Regularized Singular Vector Adaptation
@@
The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024). Nevertheless, unlike PiSSA which merge $\Sigma_{p}$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\Sigma_{p}^{\frac{1}{2}}$ and $B=\Sigma_{p}^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ in separate matrices. SORSA is defined by Eq equation 7, which is initially equivalent to the pre-trained weight $W_{0}$.
@@
We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 5.
${\mathcal{L}}_{reg}=\|U^{\top}_{p}U_{p}-I\|_{F}+\|V^{\top}_{p}V_{p}-I\|_{F}$ (8)
@@
At training step $t$, $\nabla_{W_{p,t}}{\mathcal{L}}_{train}$ denotes the gradient of ${\mathcal{L}}_{train}$ respect to $W_{p,t}$, and $\nabla_{W_{p,t}}{\mathcal{L}}_{reg}$ denotes the gradient of ${\mathcal{L}}_{reg}$ respect to the orthonormal regularizer loss $W_{p,t}$. $\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss at step $t$, respectively.
We update the SORSA as Eq equation 10 for implementation simplicity.
@@
$\eta_{d}$ is the maximum learning rate which is given from the scheduler. This implementation allows us to use only one optimizer and scheduler to deal with two different learning rates separately.
5 Gradient Analysis
In this section, we present a comprehensive mathematical analysis of the SORSA method which mainly focus on the effect of orthonormal regularization. Our investigation elucidates the key optimization properties of SORSA, providing a theoretical foundation for its advantages. We explore four critical aspects: the convexity of the regularizer, the Lipschitz continuity of the gradient, bounds on the hyperparameter $\gamma$, and the impact on the condition number of the optimization problem.
The proofs of the theorems and lemmas, along with additional mathematical details, are provided in Appendix B.
Our analysis reveals fundamental theoretical properties of SORSA, establishing its mathematical soundness and demonstrating its optimization advantages. We prove two key theorems that form the cornerstone of our theoretical framework:
Theorem 5.1.
The regularizer ${\mathcal{L}}_{reg}$ is convex.
Theorem 5.2.
The gradient of the regularizer ${\mathcal{L}}_{reg}$ is Lipschitz continuous.
These properties collectively ensure a well-behaved optimization landscape, facilitating stable and efficient training of SORSA adapters.
Building upon these foundational results, we further analyze the bounds of the hyperparameter $\gamma$, a critical factor in the performance of SORSA:
Theorem 5.3.
For convergence of gradient descent, the learning rate $\eta_{d}$ and regularization parameter $\gamma$ must satisfy:
$\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (11)
This theorem provides crucial guidance for practitioners, offering a clear criterion for selecting appropriate values of $\gamma$ to ensure convergence of the gradient descent process.
To demonstrate the superior optimization properties of SORSA, we present a novel analysis of its condition number, a critical factor in determining convergence speed and stability. Our theoretical investigation reveals a significant improvement in the condition number compared to unregularized approaches, providing a mathematical foundation for SORSA’s enhanced performance.
We begin this analysis by establishing a key lemma that bounds the effect of the orthonormal regularizer on the singular values of the weight matrix:
Lemma 5.4.
Let $W_{p}^{\text{unreg}}=U_{p}^{\text{unreg}}\Sigma_{p}^{\text{unreg}}(V_{p}^{\text{unreg}})^{\top}$ be the $W_{p}$ only training without using regularizer, and $W_{p}^{\text{reg}}=U_{p}^{\text{reg}}\Sigma_{p}^{\text{reg}}(V_{p}^{\text{reg}})^{\top}$ be the $W_{p}$ training with the regularizer. For each singular value $\sigma_{i}$, the following bound holds:
$(1-\epsilon)\sigma_{i}^{\text{unreg}}\leq\sigma_{i}^{\text{reg}}\leq(1+\epsilon)\sigma_{i}^{\text{unreg}}$ (12)
where $\epsilon$ is a small positive constant.
Lemma 5.4 provides a crucial connection between the regularizer and the singular values. Building on this result, we arrive at our main theorem regarding the condition number:
Theorem 5.5.
The orthonormal regularizer in SORSA can improve the condition number of the optimization problem over the course of training under certain conditions. Specifically:
At initialization ($t=0$):
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=1$ (13)
When training ($t>0$), while $U_{p}^{\text{unreg}}$ and $(V_{p}^{\text{unreg}})^{\top}$ are more and more deviating from orthonormal, $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ will increase, and $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will generally remain $1$ due to the regularization in orthonormality of $U_{p}^{\text{reg}}$ and $(V_{p}^{\text{reg}})^{\top}$. When $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ becomes greater than $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$, the ratio becomes strictly less than 1, indicating an improvement in the condition number.
This theorem quantifies the improvement in the condition number achieved by SORSA, offering an explanation for its fast convergence. The proof leverages the effects of the orthonormal regularization to establish a tight bound on the condition number ratio. This theorem could also shows that training with the regularizer, the distribution of $W_{p}$ will be more evenly distributed due to a smaller ratio between $\sigma_{\max}(W_{p})$ and $\sigma_{\min}(W_{p})$, which means a better training stability.
In conclusion, these theorems provide a mathematical foundation for the SORSA method. These theoretical guarantees validate the empirical success of SORSA but also provide valuable insights for future developments in PEFT methods.
@@ 6 Empirical Experiments
We conducted NLG tests on Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023). We trained the models with SORSA applied using MetaMathQA (Yu et al., 2024), and evaluated the model on GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We also trained the model on CodeFeedback Filtered Instruction (Zheng et al., 2024) dataset and evaluated it on HumanEval (Chen et al., 2021). The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024). All reported values are accuracy. See Appendix A.2 for more details and hyperparameters of the training. We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from Meng et al. (2024). For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80GB) GPU. See Table 1 for the results and Figure 3 for the loss and gradient norm comparison.
Model Method Trainable Parameters MATH GSM-8K HumanEval
Llama 2 7B Full FT 6738M 7.22 49.05 21.34
LoRA 320M 5.50 42.30 18.29
PiSSA 320M 7.44 53.07 21.95
SORSA 320M 10.36 56.03 24.39
Mistral 7B Full FT 7242M 18.60 67.02 45.12
LoRA 168M 19.68 67.70 43.90
PiSSA 168M 21.54 72.86 46.95
SORSA 168M 21.86 73.09 47.56
Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from Meng et al. (2024).
@@
Figure 3: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B. LoRA and PiSSA curves are from Meng et al. (2024).
@@ 7 Conclusion
We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis.
Our gradient analysis provides a mathematical foundation for SORSA, demonstrating its convexity, Lipschitz continuity, and the crucial role of the regularizer in improving the optimization landscape. This theoretical framework not only explains SORSA’s empirical superior performance but also offers valuable insights for future developments in adaptive learning algorithms.
@@ 8 Future Work
  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different models, in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
  • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2020) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.